OpenStudy (anonymous):

√5x+29=x+3 Use a comma to separate answers as needed.

OpenStudy (anonymous):

we can subtract 3 from both side

OpenStudy (anonymous):

1/2(57+√545)

OpenStudy (anonymous):

\[\sqrt{5}x+26=x\]

OpenStudy (anonymous):

that is what I have so far

OpenStudy (anonymous):

\[\sqrt{5}x=x-26\] square both side \[(\sqrt{5}x)^2=(x-26)^2\]

OpenStudy (anonymous):

\[5x^2=676 - 52 x + x^2\]

OpenStudy (anonymous):

(x-26)^2

OpenStudy (anonymous):

-1/2√(^3√1/2(225+√237247366593 ------------------------------- 3 2/3 ???

OpenStudy (anonymous):

I do not think that is right

OpenStudy (anonymous):

\[676-52x-x^4=0\] this is right